OpenStudy (nick2002):

Tell whether the sequence 2, -8, 32, -128 . . . is arithmetic, geometric, or neither. Find the next three terms of the sequence

jhonyy9 (jhonyy9):

than you see 2*(-4) = -8 -8*(-4) = 32 32*(-4) = -128 do you can continue it ?

OpenStudy (nick2002):

i got that but i think it arthimic

rebeccaxhawaii (rebeccaxhawaii):

nope

rebeccaxhawaii (rebeccaxhawaii):

it would not be arithmetic

OpenStudy (nick2002):

or would it be geometric because its multyplying

rebeccaxhawaii (rebeccaxhawaii):

yes look at what jhony has given you

rebeccaxhawaii (rebeccaxhawaii):

do you know how to get your next three terms?

OpenStudy (nick2002):

yes

rebeccaxhawaii (rebeccaxhawaii):

great :D good luck